# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v3

@@ A.11 How many ways should we describe a task?
A.12 Cost of elicitation
The priors elicited from GPT-3.5-turbo incur a cost, as we use OpenAI’s API. To elicit all of the priors used in the experiments in Section 3.1, we were charged \$$\$0.41$ (on average \$$\$0.068$ per dataset). However, we did not make use of the batch API and instead made each API call sequentially. All experiments in this work could be implemented using the batch API since no new API call depends on a previous API result, which may significantly reduce costs  [^5] https://openai.com/api/pricing/.
A.13 Other language models
@@
In this section, we will compare the results presented in the main paper (on GPT-3.5-turbo) to some other commonly used language models. We start by exploring the performance of some other OpenAI models  [^6] https://platform.openai.com/docs/models, before considering some of the open source Llama models  [^7] https://huggingface.co/meta-llama [15, 16, 17].
A.13.1 Other OpenAI models
@@
Figure 17: Test accuracy of the posterior distribution for different OpenAI language models. Here we show the posterior results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@
In this section we compare the posterior results of priors elicited from GPT-3.5-turbo to GPT-4-turbo which incurs a significantly higher cost, and GPT-4o-mini which provides a less expensive option.
Specifically, to elicit all priors presented here, GPT-4o-mini cost \$$\$0.32$ (on average \$$\$0.053$ per dataset), whilst GPT-4-turbo cost \$$\$7.56$ (on average \$$\$1.26$ per dataset). This is compared to the cost of \$$\$0.41$ (on average \$$\$0.068$ per dataset) for GPT-3.5-turbo (presented in A.12). As a rule, the more features a dataset contains or the longer the task description, the more expensive priors are to elicit since the number of input and output tokens is greater.
Figure 17 shows that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the synthetic and Hypothyroid prediction tasks. On the Breast Cancer, UTI, and Heart Disease tasks, the priors elicited from GPT-4-turbo allow for a greatly improved posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. In addition, on the Diabetes predictive task, GPT-4-turbo achieves similar mean squared error to GPT-3.5-turbo, with a small improvement in performance when only a few training data points are available.
@@
On the other hand, GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo in terms of their posterior performance in all tasks except one. For the synthetic, Diabetes, and Heart Disease prediction tasks, GPT-4o-mini provides prior distributions that achieve similar or worse performance to the uninformative prior. However, for the UTI and Breast Cancer predictive tasks they achieve similar posterior accuracy to GPT-3.5-turbo at a lower price point. Also, on the Hypothyroid dataset GPT-4o-mini provides prior distributions that achieve greater accuracy than those from GPT-3.5-turbo and GPT-4-turbo for all training sizes, whilst requiring significantly less cost.
A.13.2 Open source language models
@@
Figure 18: Test accuracy of the posterior distribution for Llama language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@
Here, we explore the posterior performance of prior distributions elicited from three Llama [15, 16, 17] models and one Qwen model [53, 54]:
@@
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open source. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
@@
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because these smaller models sometimes fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
@@
Figure 18 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of tested language models. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still produce the greatest posterior performance.
@@
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic and Hypothyroid tasks, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. However, for the UTI and Heart Disease datasets, both the 3B and 8B Llama models provide prior distributions that are of worse quality than even the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models are significantly smaller in size.
@@
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with similar posterior performance to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for a predictive task. However, when eliciting priors for the Diabetes and synthetic tasks, we found that Llama-3.1-70B-Instruct-BF16 provided distributions that were not competitive with GPT-3.5-turbo.
@@
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart Disease and Hypothyroid prediction, it achieves a greater posterior accuracy than all models in Figure 18, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after seeing $10$ demonstrations, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. However, on the Breast Cancer and synthetic tasks, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct provides inconsistent performance across tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.
@@
It is important, therefore, given an elicited prior distribution, to calculate the prior likelihood with some already collected data to understand whether that distribution is helpful for constructing a predictive model.

### v4

@@ A.12 How many ways should we describe a task?
A.13 Is the language model being consistent during in-context learning?
Figure 15: Distribution of the prior parameters under elicitation and those approximated from the internal model for the different predictive tasks for GPT-3.5-turbo. Here, we compare the prior distributions of the different predictive tasks, as defined by the internal model used in-context, and that which we elicit from the language model. Here, we draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
In Section 5.5, we study how consistent GPT-3.5-turbo is when we elicit a prior distribution, compared to the prior it uses in its internal predictive model when performing in-context learning.
A.14 Empirically, is the language model approximating Bayesian inference?
In this section, we present the differences between the posterior distribution extracted through maximum likelihood sampling of GPT-3.5-turbo’s in-context internal model (Section 5.4) and one that is sampled through MC methods on the extracted in-context prior distribution.
Table 4: Difference between approximated posterior and MC on the approximated prior for GPT-3.5-turbo. The energy statistic between the language model’s internal posterior extracted through MLE sampling and the posterior samples from applying Monte Carlo techniques to the language model’s internal prior. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar. Here, we show the mean and standard deviation of the energy statistic for the $5$ different training sets tested as observed data.
Heart Disease Hypothyroid Diabetes Breast Cancer
0.04 ± 0.01 0.06 ± 0.01 0.02 ± 0.00 0.01 ± 0.00
Figure 16: The different posterior samples for GPT-3.5-turbo. Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
A.15 AutoElicit with other language models
@@
In this section, we will compare the results presented in the main paper with other commonly used language models. We start by exploring the performance of some other OpenAI models, before considering some of the open-source Llama and Qwen models.
A.15.1 Other OpenAI models
@@
Figure 17: Test accuracy of the posterior distribution for some OpenAI language models. Here we show the posterior results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@
Here, we compare the results of priors elicited from GPT-3.5-turbo, DeepSeek-R1-32B-Q4 (open-source), and GPT-4-turbo (which incurs a significant cost) to GPT-4o-mini (which provides a less expensive option).
To elicit all priors presented here, GPT-4o-mini cost \$$\$0.32$ (on average \$$\$0.053$ per dataset), whilst GPT-4-turbo cost \$$\$7.56$ (on average \$$\$1.26$ per dataset) and GPT-3.5-turbo cost \$$\$0.41$ (on average \$$\$0.068$ per dataset).
@@
Figure 17 shows that GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo in terms of their posterior performance in all tasks except one. For the synthetic, Diabetes, and Heart Disease prediction tasks, GPT-4o-mini provides prior distributions that achieve similar or worse performance to the uninformative prior. However, for the UTI and Breast Cancer predictive tasks they achieve similar posterior accuracy to GPT-3.5-turbo at a lower price point. Also, on the Hypothyroid dataset GPT-4o-mini provides prior distributions that achieve greater accuracy than those from GPT-3.5-turbo and GPT-4-turbo for all training sizes, whilst requiring significantly less cost. Further, in all tasks except Diabetes progression prediction, DeepSeek-R1-32B-Q4 provides better accuracy than GPT-4o-mini.
A.15.2 Open-source language models
@@
Figure 18: Test accuracy of the posterior distribution for open-source language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@
Here, we explore the posterior performance of prior distributions elicited from three Llama (Touvron et al., 2023; MetaAI, 2024; Dubey et al., 2024) models, one Qwen model (Yang et al., 2024; Team, 2024), and one DeepSeek model (DeepSeek-AI et al., 2025) to those elicited from the OpenAI models presented in the main paper:
@@
In contrast to the OpenAI models in Appendix A.15.1, these models are publicly available and open-source. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
@@
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, GPT-4-turbo and DeepSeek-R1-32B-Q4. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.7.3.
@@
Figure 18 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of models. The 3B and 8B variants are significantly smaller in capacity to those in the main paper; justifying their inconsistent elicited priors. However, in some of the datasets they still produce good posterior performance.
@@
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic and Hypothyroid tasks, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. However, for the UTI and Heart Disease datasets, both the 3B and 8B Llama models provide prior distributions that are of worse quality than even the uninformative prior. It is possibly not surprising that Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct provided priors that were equal or worse than Llama-3.1-70B-Instruct-BF16 in all cases except for the synthetic task, since these models are significantly smaller in size.
@@
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with posterior performance similar to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo and comparable values to GPT-4 turbo or DeepSeek-R1-32B-Q4. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo and equal to DeepSeek-R1-32B-Q4. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for predictive tasks.
@@
Qwen-2.5-14B-Instruct produces priors with similar inconsistent posterior performance. For Heart Disease and Hypothyroid prediction, it achieves a competitive posterior accuracy, which is surprising given its size. For the UTI task, after seeing $10$ demonstrations, it performs comparably to Llama-3.1-70B-Instruct-BF16 and GPT-4-turbo, and better than GPT-3.5-turbo. However, on the Breast Cancer and synthetic tasks, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct provides inconsistent performance across tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.
The fact that DeepSeek-R1-32B-Q4 is an intermediate size between Qwen-2.5-14B-Instruct and Llama-3.1-70B-Instruct-BF16, and provides considerably better performance with a higher quantisation is a testament to the power of recent advances in reasoning (Huang & Chang, 2023) and test-time scaling (Jones, 2021).
@@
Therefore, it is important, given an elicited prior distribution, to calculate the prior likelihood with some already collected data to understand which prior is most appropriate for constructing a predictive model.
A.16 Choosing between priors
As seen in Appendices A.15.1 and A.15.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that DeepSeek-R1-32B-Q4, GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this may not be the case for all datasets or language models. Therefore, given prior distributions elicited from different language models, we would like to choose the best one for our predictive task.
For this, given some collected data, we can use the Bayes factor (Section 4.3) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques, such as the Bayesian Information Criterion.
Further, by using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.